# arXiv:2603.02287v1[quant-ph] 02 Mar 2026

### v1

@@ VI.2 Second-order correlation function: HBT effect
$\dfrac{d}{d\tau}\Big<\mathcal{O}_{1}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+$ $\tau^{\prime})\mathcal{O}_{2}(\tau^{\prime})\Big$>
$=$ $M\Big<\mathcal{O}_{1}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{{}^{\prime}})\mathcal{O}_{2}(\tau^{\prime})\Big$>~; (154)
@@
In providing the example of the second-order correlation function, we choose a specific set of operators. For the third component of $\bm{\mathcal{S}}$, we fix $\mathcal{O}_{1}\,=\,\sigma_{+}^{\,(B)}$ and $\mathcal{O}_{2}\,=\,\sigma_{-}^{\,(B)}$. Here, we should mention that this combination is relevant in the context of the photon bunching or the Hanbury-Brown-Twiss (HBT) [66] effect, a pivotal phenomenon in quantum optics. Our aim is to compare the outcome for our two-level system, which essentially mimics fermionic statistics, with the existing results in the literature for a multi-level system that mimics bosonic statistics.
@@
$\Big<\sigma_{+}^{\,(B)}(\tau^{\prime})\,\sigma_{+}^{\,(B)}(\tau+\tau^{\prime})\,\sigma_{-}^{\,(B)}(\tau+\tau^{\prime})\,\sigma_{-}^{\,(B)}(\tau^{\prime})\Big>_{ss}$
$=$ $\frac{1}{\Big(1+\exp[\frac{2\pi\Omega}{a}\big]\Big)^{2}}\,\bigg(\,1-e^{-\frac{\mu^{2}\Omega}{8\pi}\,\coth\left(\frac{\pi\,\Omega}{a}\right)\,\tau}\bigg)~;$ (161)
@@
$=\bigg[\frac{\mathcal{G}(-\Omega)}{\mathcal{G}_{+}(\Omega)}\bigg]^{2}=\dfrac{1}{\bigg(1\,+\,\exp[\frac{2\pi\Omega}{a}\big]\bigg)^{2}}\,=\,\big(P_{2}(\Omega)\big)^{2}~.$ (163)
At this point, we intend to mention that eq. (161) is found to be in contrast with the same found for the bosonic system. As discussed in [66], in the long time, the second order correlation function for a bosonic damped oscillator reads
@@
Once again, to verify the consistency with respect to $\big\{\sigma_{+}^{\,(B)},\sigma_{-}^{\,(B)}\big\}\,=\,1$, considering $\bm{\mathcal{S}}=\Delta=\sigma_{-}^{\,(B)}\sigma_{+}^{\,(B)}$ in eq. (154) and using eq. (116), we derived

### v2

@@ VI.2 Second-order correlation function: HBT effect
$\dfrac{d}{d\tau}\Big<\mathcal{O}_{1}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+$ $\tau^{\prime})\mathcal{O}_{2}(\tau^{\prime})\Big$>
$=$ $M\Big<\mathcal{O}_{1}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{{}^{\prime}})\mathcal{O}_{2}(\tau^{\prime})\Big$>~; (155)
@@
In providing the example of the second-order correlation function, we choose a specific set of operators. For the third component of $\boldsymbol{\mathcal{S}}$, we fix $\mathcal{O}_{1}\,=\,\sigma_{+}^{\,(B)}$ and $\mathcal{O}_{2}\,=\,\sigma_{-}^{\,(B)}$. Here, we should mention that this combination is relevant in the context of the photon bunching or the Hanbury-Brown-Twiss (HBT) [21] effect, a pivotal phenomenon in quantum optics. Our aim is to compare the outcome for our two-level system, which essentially mimics fermionic statistics, with the existing results in the literature for a multi-level system that mimics bosonic statistics.
@@
$P_{SS}$
$=$ $\Big<\sigma_{+}^{\,(B)}(\tau^{\prime})\,\sigma_{+}^{\,(B)}(\tau+\tau^{\prime})\,\sigma_{-}^{\,(B)}(\tau+\tau^{\prime})\,\sigma_{-}^{\,(B)}(\tau^{\prime})\Big>_{ss}$
$=$ $\frac{1}{\Big(1+\exp[\frac{2\pi\Omega}{a}\big]\Big)^{2}}\,\bigg(\,1-e^{-\frac{\mu^{2}\Omega}{8\pi}\,\coth\left(\frac{\pi\,\Omega}{a}\right)\,\tau}\bigg)~.$ (162)
The graphical representation of eq. (162) is shown in Figure 4. For various numerical values of the uniform acceleration $a$ (in natural units), we plot the variation of steady-state correlation function with respect to the time delay $\tau$. As evident from the graph, for a particular value of $\tau$, the correlation increases as the value of the uniform acceleration increases.
Figure 4: Plot of the two point steady-state correlation function of second order (associated with HBT effect) as we vary time delay $\tau$ for different numerical values of uniform acceleration $a$ and for $\Omega=1,\mu\,=1$.
The above expression, for $\Omega\,>0$, in zero delay ($\tau=0$) and in long delay ($\tau\rightarrow\infty$), respectively, yields
@@
$=\bigg[\frac{\mathcal{G}(-\Omega)}{\mathcal{G}_{+}(\Omega)}\bigg]^{2}=$ $\dfrac{1}{\bigg(1\,+\,\exp[\frac{2\pi\Omega}{a}\big]\bigg)^{2}}\,=\big(P_{2}(\Omega)\big)^{2}$
$\equiv\,$ $\,P_{HBT}$ (164)
Figure 5: Plot of the two point steady-state correlation function of second order (associated with HBT effect) at infinite time delay as we vary frequency $\Omega$ for different numerical values of uniform acceleration $a$.
Although Figure 5, which represents the graphical form of the above expression, is a special case of Figure 4 at $\tau\rightarrow\infty$, and it also increases with increasing uniform acceleration for a fixed value of $\Omega$, the curves exhibit different behaviour compared to those in Figure 4.
Next, we intend to mention that eq. (162) is found to be in contrast with the same found for the bosonic system. As discussed in [21], in the long time, the second order correlation function for a bosonic damped oscillator reads
@@
Once again, to verify the consistency with respect to $\big\{\sigma_{+}^{\,(B)},\sigma_{-}^{\,(B)}\big\}\,=\,1$, considering $\boldsymbol{\mathcal{S}}=\Delta=\sigma_{-}^{\,(B)}\sigma_{+}^{\,(B)}$ in eq. (155) and using eq. (117), we derived